OpenStudy (anonymous):

which of the following statements is (are) true for all postitive intergers 1) n^3+2n is divided by 3 2)5^(2n)-1 is divided by 24

OpenStudy (zzr0ck3r):

are you doing induction?

OpenStudy (anonymous):

divisible*

OpenStudy (anonymous):

yess i think

OpenStudy (zzr0ck3r):

im pretty sure 1 is true, do you need to prove it?

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

its on e2020

OpenStudy (anonymous):

i have no clue how to do it lol

OpenStudy (zzr0ck3r):

it is true for the case when n = 1 assume n^3+2n is divisible by 3 for some n in N now we show the case for n+1 (n+1)^3 +2(n+1) = (n+1)(n^2+2n+1) + 2n + 2 = (n^3+2n^2+n+n^2 + 2n + 1 + 2n +2) (n^3+3n^2+5n+3) = (n^3+2n )+ 3(n + 1+n^2) we assume (n^3+2n) is divisble by 3 and we know 3(n+1+n^2) is divisble by 3 thus by induction n^3+2n is divisible by 3 for all n in N

OpenStudy (zzr0ck3r):

I think the second one is true also, you can show it by the same method.